OpenStudy (anonymous):

Find 4x − 2y − z, if x = 2, y = 3, and z = −4

OpenStudy (anonymous):

6

OpenStudy (anonymous):

4x − 2y − z = 4(2) - 2(3) - (-4) = 8 - 6 + 4 = 6

OpenStudy (anonymous):

4(2)-2(3)-(-4) 8-6-(-4) 2+4 6

OpenStudy (anonymous):

Thank you